A 1200·kg car traveling with a velocity of 12·m/s north collides head-on with a 2400·kg truck traveling with a velocity of 8·m/s south. Immediately after the collision, the velocity of the car is 4·m/s south. Take north as the positive direction

(a) What is the velocity of the truck immediately after the collision?

(b) What impluse is delivered to the car by the truck?

(c) What impluse is delivered to the truck by the car?
Answer #1

let

m1 = 1200 kg, u1 = 12 m/s
m2 = 2400 kg, u2 = -8 m/s

after the collsion, v1 = -8 m/s

v2 = ?

a) apply conservation of momentum

m1*u1 + m2*u2 = m1*v1 + m2*v2

m2*v2 = (m1*u1 + m2*u2 - m1*v1)

v2 = (m1*u1 + m2*u2 - m1*v1)/m2

= (1200*12 + 2400*(-8) - 1200*(-8))/2400

= 2 m/s
b) Impulse delivered to the car by truck = change in momentum of the car

= m1*(v1-u1)

= 1200*(-8 - 12)

= -24000 kg.m/s or N.s

c) Impulse delivered to the truck by car = m2*(v2 - u2)

= 2400*(2 - (-8))

= 24000 kg.m/s or N.s